# Inexact Moreau Envelope Lagrangian Method for Non-Convex Constrained Optimization under Local Error Bound Conditions on Constraint Functions

### v1

@@ 6 Numerical Experiment
Figure 1: Performances of $\mathrm{iMELa}$, $\mathrm{SP}$-$\mathrm{LM}$, $\mathrm{iPPP}$ and $\mathrm{SSG}$ with static and diminishing step-sizes vs number of steps on classification problems with demographic parity fairness.
We compare the methods based on the objective value, infeasibility, stationarity, and complementary slackness (see (5), (6) and (7)) achieved by ${\mathbf{x}}^{(t)}$ and the corresponding Lagrangian multipliers. [^3] Since the SSG method does not generate a Lagrangian multiplier, we do not report its measures of stationarity and complementary slackness. While the calculations for most of these metrics are straightforward, additional details are provided for computing the stationarity metric. Observe that $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:{\mathbf{B}}{\mathbf{x}}\leq r\cdot\mathds{1}\}$, where ${\mathbf{B}}\in\mathbb{R}^{2^{d}\times d}$ is a matrix whose rows are all vectors in $\{1,-1\}^{d}$, and $\mathds{1}$ denotes the all-one vector in $\mathbb{R}^{2^{d}}$. If $\|{\mathbf{x}}\|_{1}<r$, then $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})=\{{\mathbf{0}}\}$ so that the stationarity measure in (5) reduces to $\|\mathcal{R}({\mathbf{x}})\cdot\nabla\mathcal{R}({\mathbf{x}})+\lambda\cdot\nabla\mathcal{L}({\mathbf{x}})\|$, where $\lambda\geq 0$ is the Lagrange multiplier produced by the algorithm to pair with ${\mathbf{x}}$. If $\|{\mathbf{x}}\|_{1}=r$, then by [13, Example A.5.2.6 (b)], there exists an index set of the active constraints at ${\mathbf{x}}\in\mathcal{X}$ defined by
@@
where $[{\mathbf{B}}]_{j}$ denotes the $j$-th row of ${\mathbf{B}}$. Then the normal cone to $\mathcal{X}$ at ${\mathbf{x}}$ in this case is
$\mathcal{N}_{\mathcal{X}}({\mathbf{x}})=\text{cone}\{[{\mathbf{B}}]_{j}^{\top}:j\in J({\mathbf{x}})\}=\textstyle\{\sum_{j\in J({\mathbf{x}})}a_{j}[{\mathbf{B}}_{j}]^{\top}:a_{j}\geq 0,\,j\in J({\mathbf{x}})\}.$ (93)
@@
$\min\big{\{}\big{\|}\mathcal{R}({\mathbf{x}})\cdot\nabla\mathcal{R}({\mathbf{x}})+\lambda\cdot\nabla\mathcal{L}({\mathbf{x}})+\textstyle\sum_{j\in J({\mathbf{x}})}a_{j}[{\mathbf{B}}_{j}]^{\top}\big{\|}\,\big{|}\,a_{j}\geq 0,\,j\in J({\mathbf{x}})\big{\}}.$
@@
To compare the methods in their oracle complexity, in Figure 1, we plot the four aforementioned performance metrics against the accumulated number of gradient steps performed by the APG method called within the iMEla method and the iPPP method, and the number of gradient steps performed by SP-LM and the SSG method. According to Figure 1, on all the instances studied, the iMELa method performs similarly to SP-LM and outperforms the iPPP method in all performance metrics. This is consistent with our theoretical finding that the iMELa method has an oracle complexity similar to SP-LM but better than the iPPP method. Moreover, the iMELa method and SP-LM reduce the infeasibility quickly and maintains a (nearly) feasible solution throughout the remaining iterations. In addition, we find that for all instances, the number of gradient steps performed by the APG method for solving (84) within the iMELa method is generally much smaller than that for solving (85), which is the main reason for the better empirical performance of the iMELa method than the iPPP method.
@@ 7 Conclusion
We analyze the oracle complexity of the iMELa method for finding an $\epsilon$-KKT point in smooth non-convex functional constrained optimization. Each iteration consists of approximately solving a strongly convex subproblem, followed by updates to the Lagrange multipliers and the proximal center. Our results show that the oracle complexity of the iMELa method matches the best known result in the literature, up to a logarithmic factor, under a local error bound condition, which is a strictly weaker assumption than the local LICQ condition required to achieve the best complexity guarantee.
@@
By the structure of $\mathcal{X}$ and the Slater’s condition in Assumption 3.1A and F, $\mathcal{X}^{*}$ is non-empty (see, e.g. [2, Section 3.3]). We then prove the lemma by two steps.
@@
$=$ $\leavevmode\nobreak\ \big{\langle}\boldsymbol{\xi}^{*}+\nabla f({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\big{\rangle}\leq\big{\langle}\nabla f({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\big{\rangle}\leq B_{f}D_{\mathcal{X}},$
@@
$\nabla f(\bar{{\mathbf{x}}}^{(k)})+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}^{(k)}\nabla g_{i}(\bar{{\mathbf{x}}}^{(k)})+\bar{\boldsymbol{\xi}}^{(k)}={\mathbf{0}},\leavevmode\nobreak\ g_{i}(\bar{{\mathbf{x}}}^{(k)})\leq 0,\leavevmode\nobreak\ \bar{\lambda}_{i}^{(k)}g_{i}(\bar{{\mathbf{x}}}^{(k)})=0,\leavevmode\nobreak\ \forall\,i\in[m].$
Since $\|\bar{\boldsymbol{\lambda}}^{(k)}\|\leq B_{\boldsymbol{\lambda}}$, passing to a subsequence if necessary, it holds that $\lim_{k\rightarrow\infty}\bar{\boldsymbol{\lambda}}^{(k)}=\bar{\boldsymbol{\lambda}}^{*}$ for some $\bar{\boldsymbol{\lambda}}^{*}\geq{\mathbf{0}}$. As $k\rightarrow\infty$, we conclude from the continuity of $\nabla f(\cdot)$ and $\nabla{\mathbf{g}}(\cdot)$ and the outer semi-continuity of $\mathcal{N}_{\mathcal{X}}(\cdot)$ that $\bar{\boldsymbol{\xi}}^{(k)}\rightarrow\bar{\boldsymbol{\xi}}^{*}$ for some $\bar{\boldsymbol{\xi}}^{*}\in\mathcal{N}_{\mathcal{X}}(\bar{{\mathbf{x}}}^{*})$ and
$\nabla f(\bar{{\mathbf{x}}}^{*})+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}^{*}\nabla g_{i}(\bar{{\mathbf{x}}}^{*})+\bar{\boldsymbol{\xi}}^{*}={\mathbf{0}},\leavevmode\nobreak\ g_{i}(\bar{{\mathbf{x}}}^{*})\leq 0,\leavevmode\nobreak\ \bar{\lambda}_{i}^{*}g_{i}(\bar{{\mathbf{x}}}^{*})=0,\leavevmode\nobreak\ \forall\,i\in[m],$
@@
Due to the compactness of $\mathcal{X}$ and $\mathcal{X}^{*}$ (see Lemma A.1), by passing to a subsequence if necessary, there exists some $\bar{{\mathbf{x}}}^{*}\in\mathcal{X}^{*}$ such that ${\mathbf{x}}^{(k)}\rightarrow\bar{{\mathbf{x}}}^{*}$ as $k\rightarrow\infty$. By the continuity of ${\mathbf{g}}(\cdot)$ and ${\mathbf{A}}(\cdot)-{\mathbf{b}}$, it holds that for any $i\in[m]$, if $g_{i}(\bar{{\mathbf{x}}}^{*})<0$, then $g_{i}({\mathbf{x}}^{(k)})<0$ as well when $k$ is large enough, and for any $j\in[l]$, if $[{\mathbf{A}}\bar{{\mathbf{x}}}^{*}-{\mathbf{b}}]_{j}<0$, then $[{\mathbf{A}}{\mathbf{x}}^{(k)}-{\mathbf{b}}]_{j}<0$ as well when $k$ is large enough. That means $J_{g}({\mathbf{x}}^{(k)})\subset J_{g}(\bar{\mathbf{x}}^{*})$ and $J_{A}({\mathbf{x}}^{(k)})\subset J_{A}(\bar{\mathbf{x}}^{*})$ for a large enough $k$. This contradicts to the assumption. Therefore, the conclusion holds. ∎
@@
The result follows from (12) and the Lipschitz continuity of $\nabla{\mathbf{g}}(\cdot)$. ∎
@@
Let $\delta=\min\left\{\frac{r(\delta^{\prime})}{2},\frac{\zeta}{2L\sqrt{m}}\right\}$, where $r(\cdot)$ is from Lemma A.2 and $\delta^{\prime}$ is from Lemma A.3. Suppose that ${\mathbf{x}}\in\mathcal{X}$ and $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta$ for some ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$, we then need to show that there exists $\gamma>0$ such that (11) holds for any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subset J_{A}({\mathbf{x}}^{*})$. For such $I_{g}$ and $I_{A}$, consider the minimization problem
[…]
@@
$\bar{\lambda}_{i}\cdot\left\langle\nabla g_{i}(\bar{{\mathbf{x}}}),{\mathbf{x}}-\bar{{\mathbf{x}}}\right\rangle\leq$ $\leavevmode\nobreak\ |\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})-g_{i}(\bar{{\mathbf{x}}})|+\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}$
$=$ $\leavevmode\nobreak\ |\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})|+\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}.$ (102)
For $i\in[m]\backslash I_{g}$, using (99), (100) and the convexity of $g_{i}$, we have
@@
$\bar{w}_{j}\cdot[{\mathbf{A}}({\mathbf{x}}-\bar{{\mathbf{x}}})]_{j}\leq\begin{cases}|\bar{w}_{j}|\cdot|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}|&\text{ for }j\in I_{A}\\ |\bar{w}_{j}|\cdot[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}]_{+}&\text{ for}j\in[l]\backslash I_{A}.\end{cases}$ (104)
@@
$\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\ +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\sqrt{\|\bar{\boldsymbol{\lambda}}\|^{2}+\|\bar{{\mathbf{w}}}\|^{2}}$
@@
$\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|^{2}\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\ +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\frac{2\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|}{\zeta}$

### v2

@@ 6 Numerical Experiment
Figure 1: Performances of iMELa, SP-LM, iPPP, DPALM, and SSG with static and diminishing step-sizes vs. number of steps on classification problems with demographic parity fairness.
where $\kappa$ is the slackness parameter indicating how much we are willing to increase the loss in order to reduce $|\mathcal{R}({\mathbf{x}})|$ to obtain a more fair model, and $\mathcal{X}$ is the same as that in (39). In the experiments, we set $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:\|{\mathbf{x}}\|_{1}\leq r\}$ with $r=6\|{\mathbf{a}}_{i}\|_{1}$ to satisfy Assumption 3.1A.
We now briefly justify that problem (41) satisfies Assumption 3.3 with exponent $d=1$. Specifically, we show that the constraint system satisfies the local LICQ condition at all KKT points, which implies the desired local EBC with exponent $d=1$ via Proposition 3.4. The formal statement and proof are deferred to Section E.2.
We solve problem (41) on three datasets: a9a [18], bank [35] and COMPAS [1]. Details about these datasets and how they are used to formulate (41) are provided in Section E.1. The implementation details of each method, including the choices of tuning parameters, are provided in Section E.3.
We compare the methods based on the objective value, infeasibility, stationarity, and complementary slackness (see (5), (6) and (7)) achieved by ${\mathbf{x}}^{(t)}$ and the corresponding Lagrangian multipliers. [^3] Since the SSG method does not generate a Lagrangian multiplier, we do not report its measures of stationarity and complementary slackness. While the calculations for most of these metrics are straightforward, additional details are provided for computing the stationarity metric. Observe that $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:{\mathbf{B}}{\mathbf{x}}\leq r\cdot\mathbf{1}\}$, where ${\mathbf{B}}\in\mathbb{R}^{2^{d}\times d}$ is a matrix whose rows are all vectors in $\{1,-1\}^{d}$, and $\mathbf{1}$ denotes the all-one vector in $\mathbb{R}^{2^{d}}$. If $\|{\mathbf{x}}\|_{1}<r$, then $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})=\{{\mathbf{0}}\}$ so that the stationarity measure in (5) reduces to $\|\mathcal{R}({\mathbf{x}})\cdot\nabla\mathcal{R}({\mathbf{x}})+\lambda\cdot\nabla\mathcal{L}({\mathbf{x}})\|$, where $\lambda\geq 0$ is the Lagrange multiplier produced by the algorithm to pair with $\mathbf{x}$}. If $\|{\mathbf{x}}\|_{1}=r$, then by Hiriart-Urruty and Lemaréchal [13, Example A.5.2.6 (b)], there exists an index set of the active constraints at ${\mathbf{x}}\in\mathcal{X}$ defined by
@@
where ${\mathbf{B}}]_{j}$ denotes the $j$-th row of $\mathbf{B}$}. Then the normal cone to $\mathcal{X}$ at $\mathbf{x}$} in this case is
$\mathcal{N}_{\mathcal{X}}({\mathbf{x}})=$ $\penalty 10000\ \text{cone}\{[{\mathbf{B}}]_{j}^{\top}:j\in J({\mathbf{x}})\}$
$=$ $\penalty 10000\ \textstyle\{\sum_{j\in J({\mathbf{x}})}a_{j}[{\mathbf{B}}_{j}]^{\top}:a_{j}\geq 0,\,j\in J({\mathbf{x}})\}.$ (43)
@@
$\min_{a_{j}\geq 0,\,j\in J({\mathbf{x}})}\left\|\mathcal{R}({\mathbf{x}})\cdot\nabla\mathcal{R}({\mathbf{x}})+\lambda\cdot\nabla\mathcal{L}({\mathbf{x}})+\textstyle\sum_{j\in J({\mathbf{x}})}a_{j}[{\mathbf{B}}_{j}]^{\top}\right\|.$
@@
To compare the methods in their oracle complexity, in Figure 1, we plot the four aforementioned performance metrics against the accumulated number of gradient steps performed by the APG method called within the iMELa method, the iPPP method, and DPALM, and the number of gradient steps performed by SP-LM and the SSG method. According to Figure 1, on all the instances studied, the iMELa method performs similarly to SP-LM and outperforms the iPPP method in all performance metrics. Moreover, DPALM exhibits performance comparable to iMELa and SP-LM on COMPAS, while on a9a and bank it converges more slowly and attains worse infeasibility metrics. These observations are consistent with our theoretical finding that the iMELa method has an oracle complexity similar to SP-LM and better than those of the iPPP method and DPALM. In addition, the iMELa method and SP-LM reduce the infeasibility quickly and maintain a (nearly) feasible solution throughout the remaining iterations. Finally, we find that for all instances, the number of gradient steps performed by the APG method for solving (116) within the iMELa method and for solving (118) within DPALM is generally much smaller than that for solving (117), which is the main reason for the better empirical performance of the iMELa method and DPALM than the iPPP method.
@@ 7 Conclusion
We analyze the oracle complexity of the iMELa method for finding an $\epsilon$-KKT point in smooth non-convex functional constrained optimization. Each iteration consists of approximately solving a strongly convex subproblem, followed by updates to the Lagrange multipliers and the proximal center. Our results show that the oracle complexity of the iMELa method matches the best known result in the literature, up to a logarithmic factor, under a local error bound condition with exponent $d=1$, which is a strictly weaker assumption than the local LICQ condition required to achieve the best complexity guarantee.
@@
In this section, we provide the detailed proof for Proposition 3.4, starting with a few technical lemmas.
@@
By the structure of $\mathcal{X}$ and the Slater’s condition in Assumption 3.1A and F, $\mathcal{X}^{*}$ is non-empty (see, e.g. Bertsekas [2, Sec. 3.3]). We then prove the lemma by two steps.
@@
$=$ $\penalty 10000\ \big\langle\boldsymbol{\xi}^{*}+\nabla f({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\big\rangle\leq\big\langle\nabla f({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\big\rangle\leq B_{f}D_{\mathcal{X}},$
@@
$\nabla f(\bar{{\mathbf{x}}}^{(k)})+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}^{(k)}\nabla g_{i}(\bar{{\mathbf{x}}}^{(k)})+\bar{\boldsymbol{\xi}}^{(k)}={\mathbf{0}},\penalty 10000\ g_{i}(\bar{{\mathbf{x}}}^{(k)})\leq 0,\penalty 10000\ \bar{\lambda}_{i}^{(k)}g_{i}(\bar{{\mathbf{x}}}^{(k)})=0,\penalty 10000\ \forall\,i\in[m].$
Since $\|\bar{\boldsymbol{\lambda}}^{(k)}\|\leq B_{\boldsymbol{\lambda}}$, passing to a subsequence if necessary, it holds that $\lim_{k\rightarrow\infty}\bar{\boldsymbol{\lambda}}^{(k)}=\bar{\boldsymbol{\lambda}}^{*}$ for some $\bar{\boldsymbol{\lambda}}^{*}\geq{\mathbf{0}}$. As $k\rightarrow\infty$, we conclude from the continuity of $\nabla f(\cdot)$ and $\nabla{\mathbf{g}}(\cdot)$ and the outer semi-continuity of $\mathcal{N}_{\mathcal{X}}(\cdot)$ [46, Proposition 6.6] that $\bar{\boldsymbol{\xi}}^{(k)}\rightarrow\bar{\boldsymbol{\xi}}^{*}$ for some $\bar{\boldsymbol{\xi}}^{*}\in\mathcal{N}_{\mathcal{X}}(\bar{{\mathbf{x}}}^{*})$ and
$\nabla f(\bar{{\mathbf{x}}}^{*})+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}^{*}\nabla g_{i}(\bar{{\mathbf{x}}}^{*})+\bar{\boldsymbol{\xi}}^{*}={\mathbf{0}},\penalty 10000\ g_{i}(\bar{{\mathbf{x}}}^{*})\leq 0,\penalty 10000\ \bar{\lambda}_{i}^{*}g_{i}(\bar{{\mathbf{x}}}^{*})=0,\penalty 10000\ \forall\,i\in[m],$
@@
Due to the compactness of $\mathcal{X}$ and $\mathcal{X}^{*}$ (see Lemma A.1), by passing to a subsequence if necessary, there exists some $\bar{{\mathbf{x}}}^{*}\in\mathcal{X}^{*}$ such that ${\mathbf{x}}^{(k)}\rightarrow\bar{{\mathbf{x}}}^{*}$ as $k\rightarrow\infty$. By the continuity of ${\mathbf{g}}(\cdot)$ and ${\mathbf{A}}(\cdot)-{\mathbf{b}}$, it holds that for any $i\in[m]$, if $g_{i}(\bar{{\mathbf{x}}}^{*})<0$, then $g_{i}({\mathbf{x}}^{(k)})<0$ as well when $k$ is large enough, and for any $j\in[l]$, if ${\mathbf{A}}\bar{{\mathbf{x}}}^{*}-{\mathbf{b}}]_{j}<0$, then ${\mathbf{A}}{\mathbf{x}}^{(k)}-{\mathbf{b}}]_{j}<0$ as well when $k$ is large enough. That means $J_{g}({\mathbf{x}}^{(k)})\subset J_{g}(\bar{\mathbf{x}}^{*})$ and $J_{A}({\mathbf{x}}^{(k)})\subset J_{A}(\bar{\mathbf{x}}^{*})$ for a large enough $k$. This contradicts to the assumption. Therefore, the conclusion holds. ∎
@@
Fix ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ and any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$, $I_{A}\subset J_{A}({\mathbf{x}}^{*})$. By the continuity of $\sigma_{\min}(\cdot)$ (see e.g. Horn and Johnson [15, Corollary 7.3.5 (a)]), for any $\mathbf{x}$} we have
$\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)\geq$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-\big\|\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]-\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big\|$
$=$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-\big\|\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}})-\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),\;{\mathbf{O}}\big]\big\|$
$\geq$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-\big\|\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}})-\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*})\big\|.$
Since the functions $g_{i}$, $i\in[m]$, are $L$-smooth by Assumption 3.1C, we have
$\big\|\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}})-\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*})\big\|\leq L\sqrt{|I_{g}|}\,\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq L\sqrt{m}\,\|{\mathbf{x}}-{\mathbf{x}}^{*}\|.$
Define $L_{H}:=L\sqrt{m}$ and set
$\delta^{\prime}:=\frac{\zeta}{2L_{H}}=\frac{\zeta}{2L\sqrt{m}}.$
Then for all feasible $\mathbf{x}$} satisfying $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta^{\prime}$, we obtain
$\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)\geq$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-L_{H}\|{\mathbf{x}}-{\mathbf{x}}^{*}\|$
$\geq$ $\penalty 10000\ \zeta-L_{H}\delta^{\prime}=\zeta-\tfrac{\zeta}{2}=\tfrac{\zeta}{2}.$
@@
Let $\delta=\min\left\{\frac{r(\delta^{\prime})}{2},\frac{\zeta}{2L\sqrt{m}}\right\}$, where $r(\cdot)$ is from Lemma A.2 and $\delta^{\prime}$ is from Lemma A.3. Suppose that ${\mathbf{x}}\in\mathcal{X}$ and $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta$ for some ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$, we then need to show that there exists $\gamma>0$ such that (14) holds with $d=1$ for any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subset J_{A}({\mathbf{x}}^{*})$. For such $I_{g}$ and $I_{A}$, consider the minimization problem
$\bar{{\mathbf{x}}}=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{z}}}\frac{1}{2}\|{\mathbf{z}}-{\mathbf{x}}\|^{2}\text{ s.t.}$ $\begin{array}[]{l}g_{i}({\mathbf{z}})=0,\,i\in I_{g},\,g_{i}({\mathbf{z}})\leq 0,\,i\in[m]\backslash I_{g},\\[0.43057pt] [{\mathbf{A}}{\mathbf{z}}-{\mathbf{b}}]_{j}=0,\,j\in I_{A},\,[{\mathbf{A}}{\mathbf{z}}-{\mathbf{b}}]_{j}\leq 0,\,j\in[l]\backslash I_{A}.\end{array}$ (47)
@@
$\bar{\lambda}_{i}\cdot\left\langle\nabla g_{i}(\bar{{\mathbf{x}}}),{\mathbf{x}}-\bar{{\mathbf{x}}}\right\rangle\leq$ $\penalty 10000\ |\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})-g_{i}(\bar{{\mathbf{x}}})|+\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}$
$=$ $\penalty 10000\ |\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})|+\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}.$ (52)
[…]
@@
[…]
@@
$\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\[1.93748pt] +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\sqrt{\|\bar{\boldsymbol{\lambda}}\|^{2}+\|\bar{{\mathbf{w}}}\|^{2}}$
@@
$\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|^{2}\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\[1.93748pt] +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\frac{2\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|}{\zeta}$